# Enhanced Feature Learning via Regularisation: Integrating Neural Networks and Kernel Methods

### v1

@@ 2.4.1 Kernel Perspective
Compared to multiple kernel learning, BKerNN offers notable advantages. MKL involves combining several predefined kernels, which is prone to overfitting as the number of kernels increases. Additionally, selecting the optimal kernel combination is challenging and often requires sophisticated algorithms. In contrast, BKerNN adapts the kernel through the learned weights $(w_{j})_{j\in[m]}$, making the optimisation process simpler and more efficient, as discussed in Section 3.
@@ 2.4.2 Neural Network Perspective
Our architecture can also be interpreted as a special type of neural network with one hidden layer. Recall that $\mathcal{F}_{\infty}$ is inspired by neural networks as it involves linear components $w$ followed by a non-linear part. In neural networks, this non-linear part is typically $\eta\sigma(\cdot)$, which we replaced with $g_{w}(\cdot)\in\mathcal{H}$ in our setting. The functions in $\mathcal{F}_{m}$ are expressed similarly with the number of particles $m$ equivalent to the number of neurons in the hidden layer.
As we discuss in Section 3.1, we learn the weights $(w_{j})_{j\in[m]}$ through gradient descent, while the functions $(g_{j})_{j\in[m]}$ are learned explicitly, leveraging a closed-form solution. This approach resonates with the work of Marion and Berthier (2023) and Bietti et al. (2023). Marion and Berthier (2023) examines a one-hidden layer neural network where the step-sizes for the inner layer are much smaller than those for the outer layer. They prove that the gradient flow converges to the optimum of the non-convex optimisation problem in a simple univariate setting and that the number of neurons does not need to be asymptotically large, which is a stronger result than the usual study of mean-field regimes or neural tangent kernel. Bietti et al. (2023) consider learning the link function in a non-parametric way infinitely faster than the low-rank projection subspace, which resonates with our method, although they focus Gaussian data.
We have also established that the function space $\mathcal{F}_{\infty}$ with bounded $\Omega_{0}$ is more extensive than the space of neural networks with ReLU activations in Section 2.3. In Section 3.2, we demonstrate that this enlargement is compatible with efficient optimisation.
@@ 2.5 Other Penalties
    Basic penalty: $\Omega_{{\rm basic}}(w_{1},\ldots,w_{m})=\frac{1}{2m}\sum_{j=1}^{m}\|w_{j}\|$, which we discussed in Section 2.1. In the continuous setting, it corresponds to $\frac{1}{2}\int_{\mathbb{R}^{d}}\|w\|{\rm d}\nu(w)$. This penalty, which does not target any specific pattern in the data-generating mechanism, is the one for which we provide theoretical results in Section 4. However, it does not work as well in practive as the following penalties.
@@
    Variable penalty: $\Omega_{{\rm variable}}(w_{1},\ldots,w_{m})=\frac{1}{2}\sum_{a=1}^{\min(m,d)}\big{(}\frac{1}{m}\sum_{j=1}^{m}(w_{j})_{a}^{2}\big{)}^{1/2}$, which is also equal to $\frac{1}{2\sqrt{m}}\sum_{a=1}^{\min(m,d)}\|W^{(a)}\|_{2}$. This penalty, inspired by the group Lasso (Yuan and Lin, 2006), is designed for variable selection, pushing quantities $\|W^{(a)}\|_{2}$ towards zero, thus encouraging dependence on a few variables. In the continuous setting, it corresponds to $\frac{1}{2}\sum_{a=1}^{\min(m,d)}\big{(}\int_{\mathbb{R}^{d}}|w_{a}|^{2}{\rm d}\nu(w)\big{)}^{1/2}$.
@@
    Concave variable penalty: The concave version of the penalty for variable selection, $\Omega_{{\rm concave\ variable}}(w_{1},\ldots,w_{m})=\frac{1}{2s}\sum_{a=1}^{\min(m,d)}\log\big{(}1+\frac{s}{\sqrt{m}}\|W^{(a)}\|_{2}\big{)}$, with $s\geq 0$. The appeal of the added concavity is discussed below. In the continuous setting, it corresponds to $\frac{1}{2s}\sum_{a=1}^{d}\log\big{(}1+s\int_{\mathbb{R}^{d}}(w_{a})^{2}{\rm d}\nu(w)\big{)}^{1/2}\big{)}$.
@@ 3.1 Optimisation Procedure
To solve this problem, we alternate between minimisation with respect to $\alpha$ and $c$, which is done in closed-form, and minimisation with respect to $w_{1},\ldots,w_{m}$ which is done using one step of proximal gradient descent.
@@ 3.1.3 Algorithm Pseudocode
Algorithm 1 BKerNN pseudocode
@@ 3.2 Convergence Guarantees on Optimisation Procedure
In this section, we discuss the convergence properties of the optimisation procedure. Although we do not provide a formal proof due to differentiability issues, we highlight the importance of the homogeneity of the Brownian kernel and present arguments suggesting the robustness of the optimisation process.
@@
$\min_{w_{1},\ldots,w_{m}\in\mathbb{R}^{d}}\left(\inf_{\alpha\in\mathbb{R}^{n},c\in\mathbb{R}}\frac{1}{2n}\|Y-K\alpha-c\mathds{1}_{n}\|_{2}^{2}+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2m}\sum_{j=1}^{m}\|w_{j}\|\right)$
$=\min_{w_{1},\ldots,w_{m}\in\mathbb{R}^{d}}\left(\frac{\lambda}{2}\tilde{Y}^{\top}\big{(}\tilde{K}+\lambda nI\big{)}^{-1}\tilde{Y}+\frac{\lambda}{2m}\sum_{j=1}^{m}\|w_{j}\|\right),$
@@
In both cases we minimise $F(\nu)$ (defined right below) over $\mathcal{P}(\mathbb{R}^{d})$ for the continuous case and over $\mathcal{P}_{n}(\mathbb{R}^{d})$, which is the set of probability distributions anchored at $n$ points on $\mathbb{R}^{d}$, in the $m$-particles case. $F$ is defined as
@@
While our framework aligns with Chizat and Bach (2022), we cannot directly apply their results due to the non-differentiability of $\Psi$ around zero, a common issue in such analyses. Despite this, our setup meets the crucial assumptions of convexity in $Q$ and the homogeneity of $\Psi$. See Experiment 1 in Section 5.2 for a numerical evaluation of the practical significance of the homogeneity assumption.
@@ 4.1 Gaussian Complexity
where $\mathcal{F}$ is $\mathcal{F}_{m}:=\{f\mid f(x)=c+\frac{1}{m}\sum_{j=1}^{m}g_{j}(w_{j}^{\top}x),w_{j}\in\mathcal{S}^{d-1},g_{j}\in\mathcal{H},c\in\mathbb{R}\}$ in practice for optimisation and :=$\mathcal{F}_{\infty}:=\{f\mid f(x)=c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}x)\,{\rm d}\mu(w),g_{w}\in\mathcal{H},\mu\in\mathcal{P}(\mathcal{S}^{d-1}),c\in\mathbb{R}\}$ for statistical analysis. Although we considered various penalties in Section 2.5, here we focus on $f\in\mathcal{F}_{\infty}$ with $\Omega(f)=\max(\Omega_{0}(f),c)$, where $\Omega_{0}(f)$ was defined as
$\Omega_{0}(f)=\inf_{c\in\mathbb{R},\mu\in\mathcal{P}(\mathcal{S}^{d-1}),(g_{w})_{w}\in\mathcal{H}^{\mathcal{S}^{d-1}}}\int_{\mathcal{S}^{d-1}}\|g_{w}\|_{\mathcal{H}}\,{\rm d}\mu(w),$
such that $f(\cdot)=c+\int_{\mathcal{S}^{d-1}}g_{w}(w^{\top}\cdot){\rm d}\mu(w)$ and corresponds to the basic penalty for $\Omega_{\rm weights}$. This is made possible through a well-defined mean-field limit; we leave the other penalties to future work.
@@
where $\varepsilon$ is a centred Gaussian vector with identity covariance matrix, and $\mathcal{D}_{n}:=(x_{1},\ldots,x_{n})$ is the data set consisting of i.i.d. samples drawn from the distribution of the random variable $X$. Note that it only contains the covariates, not the response.
@@
$G_{n}(\{f\in\mathcal{F}_{\infty}\mid\Omega(f)\leq D\})\leq D\left(\frac{1}{\sqrt{n}}+G_{n}\right)$
@@ 4.1.1 Dimension-Dependent Bound
Proof [Theorem 12] First, using Lemma 18, we have
@@
We bound the supremum inside of the expectation using covering balls. Let $M\in\mathbb{N}^{*}$ and $\mathcal{W}^{M}$ be such that $\forall w\in\mathcal{S}^{d-1},\exists\tilde{w}\in\mathcal{W}^{M}\subset\mathcal{S}^{d-1}$ such that $\|w-\tilde{w}\|\leq\zeta$, i.e., we have a $\zeta$-covering of the sphere with its own norm in $d$ dimensions. Fix $w\in\mathcal{S}^{d-1}$ and $\tilde{w}$ such that $\|w-\tilde{w}\|\leq\zeta$. We then have
@@
$=\prod_{i=1}^{n}\int_{\mathbb{R}}\frac{1}{\sqrt{2\pi}}e^{(2t(D_{\tilde{w}})_{i}-1)\frac{\varepsilon_{i}^{2}}{2}}=\prod_{i=1}^{n}(1-2t(D_{\tilde{w}})_{i})^{-1/2}.$
@@
Taking $t=\frac{1}{4\sum_{i=1}^{n}\|x_{i}\|^{*}}$, which fulfils the previously required conditions, we get
@@
$\leq\sqrt{\frac{\sum_{i=1}^{n}\|x_{i}\|^{*}}{n}}\left(\frac{\sqrt{4d\log(1+\frac{n}{2d})+2}}{\sqrt{n}}+\frac{\sqrt{4d}}{n}\right)$
@@ 4.1.2 Dimension-Independent Bound
$G_{n}\leq\frac{3}{n^{1/6}}\left((\log 2d)^{1/4}\mathds{1}_{*=\infty}+\mathds{1}_{*=2}\right)\left(\mathbb{E}_{\mathcal{D}_{n}}\bigg{(}\max_{i\in[n]}\left(\|X_{i}\|^{*}\right)^{2}\bigg{)}\right)^{1/4}.$
@@
Proof [Theorem 14] By applying Lemma 13, we have for any $\zeta_{1}>0$
@@
$\leq\mathbb{E}_{\varepsilon}\left(\sup_{w\in\mathcal{S}^{d-1},\|g\|_{\mathcal{H}}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}g_{\zeta_{1}}(w^{\top}x_{i})\right)+\|g-g_{\zeta_{1}}\|_{\infty}.$
@@
$\hat{G}_{n}$ $\leq\sqrt{\mathbb{E}_{\varepsilon}\left(\sup_{h\in\mathcal{F}_{1-{\rm Lip}}}\sup_{w\in\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(w^{\top}x_{i})\right)}$
$\leq\sqrt{\mathbb{E}_{\varepsilon}\left(\sup_{h\in\{h_{1},\ldots,h_{M}\}}\sup_{w\in\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i}\varepsilon_{i}h(w^{\top}x_{i})\right)+\zeta_{2}},$
by proceeding as with the covering of the unit ball of $\mathcal{H}$. We then use Slepian’s lemma (Ledoux and Talagrand, 1991, Corollary 3.14). For $h\in\{h_{1},\ldots h_{M}\},w\in\mathcal{S}^{d-1}$, let
All in all $\mathbb{E}_{\varepsilon}((X_{h,w}-X_{\tilde{h},\tilde{w}})^{2})\leq\mathbb{E}_{\varepsilon}((Y_{h,w}-Y_{\tilde{h},\tilde{w}})^{2})$ therefore we can apply Slepian’s lemma and obtain
$\leq\mathbb{E}_{\varepsilon}\left(\sup_{h\in\{h_{1},\ldots,h_{M}\},w\in\mathcal{S}^{d-1}}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}w^{\top}x_{i}+\sum_{j=1}^{m}\tilde{\varepsilon}_{j}\mathds{1}_{h=h_{j}}\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\right).$
$\leq\mathbb{E}_{\varepsilon}\left(\sup_{w\in\mathcal{S}^{d-1}}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}w^{\top}x_{i}+\sup_{h\in\{h_{1},\ldots,h_{M}\}}\sum_{j=1}^{m}\tilde{\varepsilon}_{j}\mathds{1}_{h=h_{j}}\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\right)$
$\leq\mathbb{E}_{\varepsilon}\left(\sup_{w\in\mathcal{S}^{d-1}}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}w^{\top}x_{i}+\sup_{j\in[M]}\tilde{\varepsilon}_{j}\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\right)$
$\leq\mathbb{E}_{\varepsilon}\left(\bigg{\|}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}x_{i}\bigg{\|}^{*}+\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\sqrt{2\log M}\right).$ (13)
We consider each term of Equation (4.1.2) separately, while also taking expectation with regards to the data set. For the second term, using the bound on $M$ (Luxburg and Bousquet, 2004, Theorem 17) and basic inequalities to simplify the term, we have
@@
$G_{n}^{2}$ $\leq\mathbb{E}_{\varepsilon,\mathcal{D}_{n}}\left(\bigg{\|}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}x_{i}\bigg{\|}^{*}\right)+\frac{8}{\sqrt{n}}\frac{1}{\sqrt{\zeta_{2}}}\left(\mathbb{E}_{\mathcal{D}_{n}}\left(\max_{i\in[n]}(\|x_{i}\|^{*})^{2}\right)\right)^{3/4}+\zeta_{2}$
@@
$G_{n}^{2}$ $\leq\left(\frac{2}{\sqrt{n}}\sqrt{\log 2d}\mathds{1}_{*=\infty}+\frac{\sqrt{2}}{\sqrt{n}}\mathds{1}_{*=2}\right)\sqrt{\mathbb{E}_{X}((\|X\|^{*})^{2})}+2\frac{4}{n^{1/3}}\sqrt{\mathbb{E}_{\mathcal{D}_{n}}\left(\max_{i\in[n]}(\|x_{i}\|^{*})^{2}\right)}$
@@
$G_{n}$ $\leq\left(\sqrt{\log 2d}\mathds{1}_{*=\infty}+\mathds{1}_{*=2}\right)^{1/4}\frac{3}{n^{1/6}}\left(\mathbb{E}_{\mathcal{D}_{n}}\left(\max_{i\in[n]}(\|x_{i}\|^{*})^{2}\right)\right)^{1/4},$

### v2

@@ 2.4.1 Kernel Perspective
Compared to multiple kernel learning, BKerNN offers notable advantages. Multiple kernel learning (MKL) involves combining several predefined kernels, which is prone to overfitting as the number of kernels increases. Additionally, selecting the optimal kernel combination is challenging and often requires sophisticated algorithms. In contrast, BKerNN adapts the kernel through the learned weights $(w_{j})_{j\in[m]}$, making the optimisation process simpler and more efficient, as discussed in Section 3.
@@ 2.4.2 Neural Network Perspective
Our architecture can also be interpreted as a special type of neural network with one hidden layer. Recall that $\mathcal{F}_{\infty}$ is inspired by neural networks as it involves linear components $w$ followed by a non-linear part. In neural networks, this non-linear part is typically $\eta\sigma(\cdot)$, which we replaced with $g(\cdot)\in\mathcal{H}$ in our setting. The functions in $\mathcal{F}_{m}$ are expressed similarly with the number of particles $m$ equivalent to the number of neurons in the hidden layer.
As we discuss in Section 3.1, we learn the weights $(w_{j})_{j\in[m]}$ through gradient descent, while the functions $(g_{j})_{j\in[m]}$ are learned explicitly, leveraging a closed-form solution. This approach resonates with the work of Marion and Berthier (2023) and Bietti et al. (2023). Marion and Berthier (2023) examine a one-hidden layer neural network where the step-sizes for the inner layer are much smaller than those for the outer layer. They prove that the gradient flow converges to the optimum of the non-convex optimisation problem in a simple univariate setting and that the number of neurons does not need to be asymptotically large, which is a stronger result than the usual study of mean-field regimes or neural tangent kernel. Bietti et al. (2023) consider learning the link function in a non-parametric way infinitely faster than the low-rank projection subspace, which resonates with our method, although they focus on Gaussian data.
We have also established that the function space $\mathcal{F}_{\infty}$ is more extensive than the space of neural networks with ReLU activations in Section 2.3. In Section 3.2, we demonstrate that this enlargement is compatible with efficient optimisation.
@@ 2.5 Other Penalties
    Basic penalty: $\Omega_{{\rm basic}}(w_{1},\ldots,w_{m})=\frac{1}{2m}\sum_{j=1}^{m}\|w_{j}\|$, which we discussed in Section 2.1. In the continuous setting, it corresponds to $\frac{1}{2}\int_{\mathbb{R}^{d}}\|w\|{\rm d}\nu(w)$. This penalty, which does not target any specific pattern in the data-generating mechanism, is the one for which we provide theoretical results in Section 4. However, it does not work as well in practice as the following penalties.
@@
    Variable penalty: $\Omega_{{\rm variable}}(w_{1},\ldots,w_{m})=\frac{1}{2}\sum_{a=1}^{{\color[rgb]{0,0,0}d}}\big{(}\frac{1}{m}\sum_{j=1}^{m}(w_{j})_{a}^{2}\big{)}^{1/2}$, which is also equal to $\frac{1}{2\sqrt{m}}\sum_{a=1}^{\min(m,d)}\|W^{(a)}\|_{2}$. This penalty, inspired by the group Lasso (Yuan and Lin, 2006), is designed for variable selection, pushing quantities $\|W^{(a)}\|_{2}$ towards zero, thus encouraging dependence on a few variables. In the continuous setting, it corresponds to $\frac{1}{2}\sum_{a=1}^{\min(m,d)}\big{(}\int_{\mathbb{R}^{d}}|w_{a}|^{2}{\rm d}\nu(w)\big{)}^{1/2}$.
@@
    Concave variable penalty: The concave version of the penalty for variable selection, $\Omega_{{\rm concave\ variable}}(w_{1},\ldots,w_{m})=\frac{1}{2s}\sum_{a=1}^{{\color[rgb]{0,0,0}d}}\log\big{(}1+\frac{s}{\sqrt{m}}\|W^{(a)}\|_{2}\big{)}$, with $s\geq 0$. The appeal of the added concavity is discussed below. In the continuous setting, it corresponds to $\frac{1}{2s}\sum_{a=1}^{d}\log\big{(}1+s\int_{\mathbb{R}^{d}}(w_{a})^{2}{\rm d}\nu(w)\big{)}^{1/2}\big{)}$.
@@
Note that the link with infinite-dimensional normed function spaces and the theoretical analysis of Section 4. only applies to the first “basic” penalty. Extending the analysis of concave penalties from finite dimension (Fan and Li, 2001; Zhang, 2010) to infinite dimension remains an open problem.
@@ 3.1 Optimisation Procedure
To solve this problem, we alternate between exact minimisation with respect to $\alpha$ and $c$, which is done in closed-form, and one improvement step for the minimisation with respect to $w_{1},\ldots,w_{m}$ which is done using one step of proximal gradient descent. This corresponds to minimizing with respect to the $w_{1},\ldots,w_{m}$ the infimum with respect to $c$ and $\alpha$.
@@ 3.1.3 Algorithm Pseudocode
Algorithm 1 BKerNN pseudocode.
@@ 3.2 Convergence Guarantees on Optimisation Procedure
In this section, we discuss the convergence properties of the optimisation procedure. Although we do not provide a formal proof due to differentiability issues, we highlight the importance of the homogeneity of the Brownian kernel and present arguments suggesting the robustness of the optimisation process. However, no formal results are presented because of the lack of differentiability.
@@
$\min_{w_{1},\ldots,w_{m}\in\mathbb{R}^{d}}\bigg{\{}\inf_{\alpha\in\mathbb{R}^{n},c\in\mathbb{R}}\frac{1}{2n}\|Y-K\alpha-c\mathds{1}_{n}\|_{2}^{2}+\frac{\lambda}{2}\alpha^{\top}K\alpha+\frac{\lambda}{2m}\sum_{j=1}^{m}\|w_{j}\|\bigg{\}}$
$=\min_{w_{1},\ldots,w_{m}\in\mathbb{R}^{d}}\bigg{\{}\frac{\lambda}{2}\tilde{Y}^{\top}\big{(}\tilde{K}+\lambda nI\big{)}^{-1}\tilde{Y}+\frac{\lambda}{2m}\sum_{j=1}^{m}\|w_{j}\|\bigg{\}},$
@@
In both cases we minimise $F(\nu)$ (defined right below) over $\mathcal{P}(\mathbb{R}^{d})$ for the continuous case and over $\mathcal{P}_{m}(\mathbb{R}^{d})$, which is the set of probability distributions anchored at $m$ points on $\mathbb{R}^{d}$, in the $m$-particles case. $F$ is defined as
@@
While our framework aligns with that of Chizat and Bach (2022), we cannot directly apply their results due to the non-differentiability of $\Psi$ around zero, a common issue in such analyses. Despite this, our setup meets the crucial assumptions of convexity in $Q$ and the homogeneity of $\Psi$. See Experiment 1 in Section 5.2 for a numerical evaluation of the practical significance of the homogeneity assumption. Obtaining formal results despite the lack of differentiability remains an open problem.
@@ 4.1 Gaussian Complexity
where $\mathcal{F}$ is $\mathcal{F}_{m}:=\{f\mid f(x)=c+\frac{1}{m}\sum_{j=1}^{m}g_{j}(w_{j}^{\top}x),w_{j}\in\mathcal{S}^{d-1},g_{j}\in\mathcal{H},c\in\mathbb{R}\}$ in practice for optimisation and :=$\mathcal{F}_{\infty}:=\{f\mid f(x)=c+\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}x)\,{\rm d}\mu(g,w),\mu\in\mathcal{P}(\mathcal{H}\times\mathcal{S}^{d-1}),c\in\mathbb{R},{\color[rgb]{0,0,0}\Omega_{0}(f)<\infty}\}$ for statistical analysis. Although we considered various penalties in Section 2.5, here we focus on $f\in\mathcal{F}_{\infty}$ with $\Omega(f)=\max(\Omega_{0}(f),c)$, where $\Omega_{0}(f)$ was defined as
$\Omega_{0}(f)=\inf_{{\color[rgb]{0,0,0}\mu\in\mathcal{P}(\mathcal{H}\times\mathcal{S}^{d-1})}}{\color[rgb]{0,0,0}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}\|g\|_{\mathcal{H}}\,{\rm d}\mu(g,w),}$
such that $f(\cdot)=c+{\color[rgb]{0,0,0}\int_{\mathcal{H}\times\mathcal{S}^{d-1}}g(w^{\top}\cdot){\rm d}\mu(g,w)}$ and corresponds to the basic penalty for $\Omega_{\rm weights}$. This is made possible through a well-defined mean-field limit; we leave the other penalties for future work.
@@
with $\varepsilon$ a centred Gaussian vector with identity covariance matrix, and $\mathcal{D}_{n}:=(x_{1},\ldots,x_{n})$ is the data set consisting of i.i.d. samples drawn from the distribution of the random variable $X$. Note that it only contains the covariates, not the response.
@@
$G_{n}(\{f\in\mathcal{F}_{\infty}\mid\Omega(f)\leq D\})\leq D\left(\frac{1}{\sqrt{n}}+G_{n}\right),$
@@ 4.1.1 Dimension-Dependent Bound
Proof [Proof of Theorem 12] First, recall the definition of $G_{n}$
$G_{n}=\mathbb{E}_{\varepsilon,\mathcal{D}_{n}}\bigg{(}\sup_{f=g(w^{\top}\cdot),\|g\|_{\mathcal{H}}\leq 1,w\in\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}f(x_{i})\bigg{)},$
which we can rewrite as
$G_{n}=\mathbb{E}_{\varepsilon,\mathcal{D}_{n}}\bigg{(}\sup_{\|g\|_{\mathcal{H}}\leq 1,w\in\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}g(w^{\top}x_{i})\bigg{)},$
and finally by splitting the supremum as everything is finite
$G_{n}=\mathbb{E}_{\varepsilon,\mathcal{D}_{n}}\bigg{(}\sup_{w\in\mathcal{S}^{d-1}}\sup_{\|g\|_{\mathcal{H}}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}g(w^{\top}x_{i})\bigg{)}.$
Using Lemma 18, we then have
@@
We bound the supremum inside of the expectation using covering balls. Let $M\in\mathbb{N}^{*}$ and $\mathcal{W}^{M}$ be such that $\forall w\in\mathcal{S}^{d-1},\exists\tilde{w}\in\mathcal{W}^{M}\subset\mathcal{S}^{d-1}$ such that $\|w-\tilde{w}\|\leq\zeta$, i.e., we have a $\zeta$-covering of the sphere with its own norm in $d$ dimensions, with $M$ the covering number. Fix $w\in\mathcal{S}^{d-1}$ and $\tilde{w}$ such that $\|w-\tilde{w}\|\leq\zeta$. We then have
@@
$=\prod_{i=1}^{n}\int_{\mathbb{R}}\frac{1}{\sqrt{2\pi}}e^{(2t(D_{\tilde{w}})_{i}-1)\frac{\varepsilon_{i}^{2}}{2}}\,{\rm d}\varepsilon_{i}=\prod_{i=1}^{n}(1-2t(D_{\tilde{w}})_{i})^{-1/2}.$
@@
Let us first check that $t=\frac{1}{4\sum_{i=1}^{n}\|x_{i}\|^{*}}$, fulfils the previously required conditions. This is equivalent to showing that we have for any $\tilde{w}\in\mathcal{W}^{M},\max_{i\in[n]}(D_{\tilde{w}})_{i}\leq\sum_{i=1}^{n}\|x_{i}\|^{*}$. First, $\max_{i\in[n]}(D_{\tilde{w}})_{i}$ is the largest eigenvalue of $K^{(\tilde{w})}$, hence it is smaller than the trace of $K^{(\tilde{w})}$. We also know that this trace is equal to $\sum_{i=1}^{n}|\tilde{w}^{\top}x_{i}|$. Now since $\|\tilde{w}\|=1$, this means that $\sum_{i=1}^{n}|\tilde{w}^{\top}x_{i}|\leq\sum_{i=1}^{n}\|x_{i}\|^{*}$. Thus by taking this $t$, we get
@@
$\leq\sqrt{\frac{\sum_{i=1}^{n}\|x_{i}\|^{*}}{n}}\left(\frac{\sqrt{4d\log(1+\frac{n}{2d})+2}}{\sqrt{n}}+{\color[rgb]{0,0,0}\sqrt{\frac{4d}{n}}}\right)$
@@ 4.1.2 Dimension-Independent Bound
$G_{n}\leq\frac{{\color[rgb]{0,0,0}6}}{n^{1/6}}\left((\log 2d)^{1/4}\mathds{1}_{*=\infty}+\mathds{1}_{*=2}\right)\left(\mathbb{E}_{\mathcal{D}_{n}}\bigg{(}\max_{i\in[n]}\left(\|X_{i}\|^{*}\right)^{2}\bigg{)}\right)^{1/4}.$
@@
Proof [Proof of Theorem 14] By applying Lemma 13, we have for any $\zeta_{1}>0$
@@
$\leq\mathbb{E}_{\varepsilon}{\color[rgb]{0,0,0}\left(\sup_{w\in\mathcal{S}^{d-1},\|g\|_{\mathcal{H}}\leq 1}\left(\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}g_{\zeta_{1}}(w^{\top}x_{i})+\|g-g_{\zeta_{1}}\|_{\infty}\right)\right)}.$
@@
$\hat{G}_{n}$ $\leq{\color[rgb]{0,0,0}2}\sqrt{\mathbb{E}_{\varepsilon}\left(\sup_{h\in\mathcal{F}_{1-{\rm Lip}}}\sup_{w\in\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(w^{\top}x_{i})\right)}$
$\leq{\color[rgb]{0,0,0}2}\sqrt{\mathbb{E}_{\varepsilon}\left(\sup_{h\in\{h_{1},\ldots,h_{M}\}}\sup_{w\in\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i}\varepsilon_{i}h(w^{\top}x_{i})\right)+\zeta_{2}},$
by proceeding as with the covering of the unit ball of $\mathcal{H}$.
We then use Lemma 19 to bound the expectation on the supremum of the finite set of Lipschitz functions, which is inspired by Bartlett and Mendelson (2002). This yields
$\mathbb{E}_{\varepsilon}$ $\bigg{(}\sup_{h\in\{h_{1},\ldots,h_{M}\},w\in\mathcal{S}^{d-1}}\frac{1}{n}\sum_{i=1}^{n}\varepsilon_{i}h(w^{\top}x_{i})\bigg{)}$
${\color[rgb]{0,0,0}\leq\mathbb{E}_{\varepsilon}\left(\bigg{\|}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}x_{i}\bigg{\|}^{*}+\sqrt{8\frac{\sum_{i=1}^{n}(\|x_{i}\|^{*})^{2}}{n^{2}}}\sqrt{2\log M}\right).}$ (13)
We then consider each term of Equation (4.1.2) separately, while also taking expectation with regards to the data set. For the second term, using the bound on $M$ (Luxburg and Bousquet, 2004, Theorem 17) and basic inequalities to simplify the term, we have
@@
${\color[rgb]{0,0,0}\frac{G_{n}^{2}}{4}}$ $\leq\mathbb{E}_{\varepsilon,\mathcal{D}_{n}}\left(\bigg{\|}\frac{\sqrt{2}}{n}\sum_{i=1}^{n}\varepsilon_{i}x_{i}\bigg{\|}^{*}\right)+\frac{8}{\sqrt{n}}\frac{1}{\sqrt{\zeta_{2}}}\left(\mathbb{E}_{\mathcal{D}_{n}}\left(\max_{i\in[n]}(\|x_{i}\|^{*})^{2}\right)\right)^{3/4}+\zeta_{2}$
@@
${\color[rgb]{0,0,0}\frac{G_{n}^{2}}{4}}$ $\leq\left(\frac{2}{\sqrt{n}}\sqrt{\log 2d}\mathds{1}_{*=\infty}+\frac{\sqrt{2}}{\sqrt{n}}\mathds{1}_{*=2}\right)\sqrt{\mathbb{E}_{X}((\|X\|^{*})^{2})}+2\frac{4}{n^{1/3}}\sqrt{\mathbb{E}_{\mathcal{D}_{n}}\left(\max_{i\in[n]}(\|x_{i}\|^{*})^{2}\right)}$
@@
$G_{n}$ $\leq\left(\sqrt{\log 2d}\mathds{1}_{*=\infty}+\mathds{1}_{*=2}\right)^{1/4}\frac{{\color[rgb]{0,0,0}6}}{n^{1/6}}\left(\mathbb{E}_{\mathcal{D}_{n}}\left(\max_{i\in[n]}(\|x_{i}\|^{*})^{2}\right)\right)^{1/4},$